A merry-go-round has a mass of 1440 kg and a radius of 7.50 m. how much net work is required to accelerate it from rest to a rotation rate of 1.00 revolution per 7.00 s? Assume it is a solid cylinder.
Answer #1

work = change in KE

W final = 2pi/7 rad/s OR 2*3.14)/7 = 0.897 rad/s

=> work = change in KE = 1/2*IWfinal^2 -0 = 1/2*(MR^2/2)*W^2 = 1/2*1440*(7.50)^2*(0.897)^2/2 = 16293.3 Joule